OpenStudy (anonymous):

solve equation using sqrt property. express radicals in simplest form: 4-3x=33/(4-3x)

hero (hero):

\[4-3x = \frac{33}{4-3x}\]

hero (hero):

\[(4-3x)^2 = 33\]

hero (hero):

\[4-3x = \sqrt{33}\]

hero (hero):

You can solve for x from there

OpenStudy (anonymous):

so... (sqrt(33)-4)/-3 does that look right?

hero (hero):

Yes

OpenStudy (anonymous):

so since I need a positive and a negative answer, will the second one be: (sqrt(33)+4)/-3 or does the negative 3 turn into a positive 3?

hero (hero):

put +/- sqrt{33}

OpenStudy (anonymous):

sry, I don't quite follow...

hero (hero):

\[x = \frac{\pm \sqrt{33}-4}{-3}\]

OpenStudy (anonymous):

nevermind, I got it. the 3 wasn't negative on either side. oh well.

OpenStudy (anonymous):

thank you though!!

hero (hero):

What do you mean it wasn't negative on either side?

OpenStudy (anonymous):

answer was: (4-sqrt33)/3, (4+sqrt33)/3

hero (hero):

It's the same answer

OpenStudy (anonymous):

I guess the system is picky

hero (hero):

If I had just posted what x was, you may have gotten it. I didn't know what form they wanted the answer in. Online tests are stupid.

OpenStudy (anonymous):

lol